element 35.3g of M reacts with nitrogen to form 43.5g of M3N2. whats the molar mass and name of the element M?




1
Expert's answer
2022-03-24T15:21:01-0400

1. The molar mass of the element M is 40 g/mol

2. The name of the element is calcium (Ca)


Description of mole

The mole an element is related to it's mass and molar mass according to the following equation:

Mole = mass / molar mass

How to determine the mass of nitrogen

  • Mass of M = 35.3 g
  • Mass of compound = 43.5 g
  • Mass of nitrogen =?

Mass of nitrogen = 43.5 – 35.3

Mass of nitrogen = 8.2 g

1. How to determine the molar mass of M

  • Mass of nitrogen = 8.2 g
  • Molar mass of nitrogen = 14 g/mol
  • Formula of compound = M₃N₂
  • Mole ratio of M to nitrogen = 3 : 2
  • Mass of M = 35.3 g
  • Molar mass of M = y =?

Mole of M : Mole of nitrogen = 3 : 2

35.3 / y ÷ 8.2 / 14 = 3 : 2

(35.3 / y) × (14 / 8.2) = 3 / 2

494.2 / 8.2y = 3 / 2

Cross multiply

8.2y × 3 = 494.2 × 2

24.6y = 988.4

Divide both side by 24.6

y = 988.4 / 24.6

y ≈ 40 g/mol

Thus, the molar mass of the element is 40 g/mol

2. How to determine the name of the element

  • Molar mass of element = 40 g/mol
  • Name of element =?

From the periodic table, the element having a mass number of 40 is Argon and Calcium. Since argon will not react with nitrogen to produce Ar₃N₂.

Therefore, we can conclude that the element is calcium since it can react with nitrogen as illustrated by the equation

3Ca + N₂ —> Ca₃N₂